OpenStudy (scarlettfarra2000):

Last winter Armand had 5/6 of a row of stacked logs. At the end of the winter he had 8/15 of the same row left. How much wood did he burn over the winter?

OpenStudy (scarlettfarra2000):

Choices A) \[1 \frac{ 9 }{ 16 }\] B) \[1 \frac{ 11 }{ 30 }\] C) \[\frac{ 4 }{ 9 }\] D) \[\frac{ 3 }{ 10 }\]